A box contains 4 red, 5 green, and 6 white balls. two balls are drawn at random from the box. what is the probability that the balls drawn include only red and/or green??
1
Expert's answer
2012-05-25T10:29:38-0400
It doesnt matter for us what colour does ball that we draw has, as far it's not
white. So, when we draw first time, the probability of drawing not white ball is
(4+5)/(4+5+6)=0.6. Next, knowing that we draw not white ball drawing out another
not white is 8/(8+6)=0.571. The answer for our question is a product of these
two numbers: 0.6*0.571=0.343